OpenStudy (anonymous):

If the measure of angle T is (123-5x)º, what is the measure of its complement?

OpenStudy (anonymous):

(123-5x)º + compliment = 90º Solve for the compliment.

OpenStudy (anonymous):

How do you do that?

OpenStudy (anonymous):

You subtract (123-5x)º from both sides. Then you simplify.

OpenStudy (anonymous):

Can u show me? please :)

OpenStudy (anonymous):

I'll write \(c\) for compliment. \[(123-5x)º + c= 90º\] \[ c= 90º -(123-5x)º \] \[ c= (5x + 90- 123)º \] \[ c= (5x - 33)º \]

OpenStudy (anonymous):

Thanks!!! that makes sense now